# Float Self-Tagging

### v1

@@ 8. Conclusion
This paper has presented a new approach to object tagging that allows attaching type information to 64-bit values at low cost. The core idea is to superimpose tags to bits that appear with high probability in practice. This approach is especially useful to represent IEEE754 floating-point numbers, whose encoding requires all bits of a 64-bit machine word.
In practice, self-tagging unboxes all floats, reducing the strain on the garbage collector and improving performance. Contrary to NaN-tagging, which unboxes all floats, self-tagging adds no overhead to the implementation of other types. Since it does not rely on the specificity of the IEEE754 NaN encoding, self-tagging is highly portable, including to 32-bit architectures.
It is also straightforward to retrofit in implementations that already use object tagging. The only required changes are to reserve one or more tags to self-tagging and implement the encoding and decoding of the target type. The rest of the runtime stays unchanged.
Self-tagging is a good candidate for languages such as JavaScript, that rely heavily on floats, but has potential applications in any polymorphic language.

### v2

@@ 9. Conclusion
Avoiding heap allocation of floating point numbers has always been a major concern for dynamic and polymorphic languages. This paper presents and evaluates the popular techniques deployed to that end and it presents a new approach, coined self-tagging, that allows attaching type information to 64-bit values at low cost. The core idea is to superimpose tags to bits that appear with high probability in practice. This approach is especially effective to represent IEEE754 floating-point numbers, whose encoding requires all bits of a 64-bit machine word.
Float self-tagging avoids heap-allocation of frequently occurring float values, reducing the strain on the garbage collector and improving performance. Contrary to NaN-boxing and NuN-boxing, which are designed to optimize the handling of floats, self-tagging offers a different trade off that adds no overhead to the handling of other types. Since it does not rely on the specificity of the IEEE754 NaN encoding, self-tagging is highly portable, including to 32-bit architectures.
It is also straightforward to retrofit in implementations that already use object tagging. The only required changes are to reserve one or more tags to float self-tagging and implement the encoding and decoding of the target type. The rest of the runtime stays unchanged.
As shown by the experimental evidence presented in the paper, float self-tagging is a good candidate for languages such as JavaScript, whose specification relies heavily on floats, but has potential applications in any polymorphic language.
Given the purely bit-shuffling nature of float self-tagging boxing and unboxing, it may also be a good candidate for an essentially zero-cost implementation at the hardware level when floats are moved between float and integer registers.